OpenStudy (anonymous):

a triangular lot has a perimeter of 1900ft one side is 200ft longer than the shortest side while the third side is 500ft longer than the shortest side find the lengths of all three sides

OpenStudy (anonymous):

let shortest side be = m then other side is = m+200 third side is = m+500 ATQ m + m+ 200 + m + 500 = 1900 3m + 700 = 1900

OpenStudy (anonymous):

assign variables shortest to largest a -> b - > c a+b+c=1900 a+200=b a+500=c

OpenStudy (anonymous):

plug n chug

OpenStudy (anonymous):

m = (1900-700)/3 = 1200/3 = 400 so shortest side is = m = 400m second side = m+200 = 600m third side = m+500 = 900m

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

pls click GOOD ANSWER button for us :))